OpenStudy (anonymous):

a sound pulse emitted underwater reflects off a school of fish and is detected at the same place 0.01 s later. how far away are the fish?

OpenStudy (abhisar):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

According to the question the sound wave takes 0.01 seconds to reach the school of fish and then come back to the place from where it was emitted. 

OpenStudy (abhisar):

Let us suppose the distance from the place of origin of sound to the school of fish be d, then sound took 0.01 seconds to travel a distance 2d. And the speed of sound is 300 m/s

OpenStudy (anonymous):

Careful, underwater this is not the case.

OpenStudy (abhisar):

Wait actually underwater is the case. So the speed of sound will be 1497 m/s @LolWolf

OpenStudy (abhisar):

Since \(\huge\ Time=\frac{Distance}{speed} \\ \huge\ 0.01=\frac{2d}{1497}\)

OpenStudy (abhisar):

Now by solving this we will get d=7.48\(\approx\)7.5m

OpenStudy (anonymous):

Also, the result slightly differs depending on whether it is fresh water or seawater. In seawater the speed of sound is slightly higher, 1560 m/s, so \[d = \frac{ 1560 m/s \times 0.01s }{ 2 } = 7.8 m\]

OpenStudy (abhisar):

@Fusky \(\huge\checkmark\)